OpenStudy (s.gracex):

using the slope-intercept form of a line, find the equation of the line with the slope 2/3 and y intercept of -2.

OpenStudy (texaschic101):

slope intercept form is y = mx + b where m is your slope and b is your y intercept. It would be written like this : y = 2/3x + (-2)....or y = 2/3x - 2

OpenStudy (s.gracex):

Thanks!

OpenStudy (texaschic101):

anytime :)